OpenStudy (anonymous):

(x – 4)2 + (y – 2)2 = 25 how do i solve this step by step?

OpenStudy (anonymous):

2x-8+2y-4=25

OpenStudy (anonymous):

2x+2y=37

OpenStudy (anonymous):

x+y=37/2

OpenStudy (anonymous):

y=37/2-x

OpenStudy (amoodarya):

(x – 4)2 + (y – 2)2 = 25 solve for x (x – 4)2 =- (y – 2)2 + 25 (x – 4)2 =25- (y – 2)2 now sqrt of both |x-4|=sqrt(25- (y – 2)2 ) x-4=+-sqrt(25- (y – 2)2 ) x=4+-sqrt(25- (y – 2)2 ) like that for y